OpenStudy (anonymous):

5x - 2y = -1 x + 4y = 35 solve using the substitution method

OpenStudy (anonymous):

also can you solve using the addition method?

OpenStudy (anonymous):

5x-2y=-x+4y 6x-6y=0 x=y

OpenStudy (angela210793):

5x - 2y = -1 x + 4y = 35 ---> x=35-4y---> 5(35-4y)-2y=-1 175-22y=-1 -22y=-176 y=8 x=35-4*8=35-32=4

OpenStudy (angela210793):

5x - 2y = -1 (*-2) 10x-4y=-2 add both sides 11x=33-->x=3 x + 4y = 35 x + 4y = 35 4y=35-3 y=8

OpenStudy (angela210793):

Solution is (3;8)

OpenStudy (anonymous):

so the first one is the substitution and the y=8 is the addition one?

OpenStudy (angela210793):

1st post is substitution and 2nd post is addition

OpenStudy (anonymous):

alright, i'm looking them over to see if i understand- thank YOU!

OpenStudy (angela210793):

tell me if u don't understand anything :D